Grade 12th passGeneral Physics

A mass M is kept hanging by rod of length L. What tangential velocity must be given to it so that it can just reach the top of vertical circle?

Bhausaheb
7 Years agoGrade 12th pass

1 Answer

Khimraj
7 Years ago

Suppose the mass of rod is ‘m’ then MI of the system is

I1=MI of the rod about the axis passing through point of suspension of the rod +MI of the mass ‘M’ about the same axis

=[(mL^2)/3+ML^2]

If V is the velocity of mass ‘M’required for the given condition then angular velocity'w’ at the lowest point is

w=V/L

Thus KE 'K’ of the system is

K=(1/2)[(mL^2)/3+ML^2] (w^2)=(1/2)[m/3+M] V^2

[Since V=L w]

Position of CM of the system'y’ below the point of suspension of the rod is given as

y=[mL/2+ML]/(m+M)=L [m/2+M]/(m+M)

At the highest point potential energy ‘U’[K is zero] is

U=(m +M)g L[m+2M]/(m+M)

=(m+2M) gL

Applying conservation law of energy we get

( 1/2)[m/3+M] V^2=(m+2M) j g L

Or [m+3M] V^2=6 (m+2M) g L

Or. V^2=6 (m+2M) gL/(m+3M)

Or. V=[6 (m+2M)gL /(m+3M)]^(1/2)

If mass of the rod ‘m’ is negligible then

V=[4gL]^(1/2)